OpenStudy (anonymous):

What are the x– and y–intercepts of y = 2x^2 – 10x + 8? Answer A. x–intercepts: (1, 0) and (4, 0) y–intercept: (0, 8) B. x–intercepts: (1, 0) and (4, 0) y–intercept: (0, 4) C. x–intercepts: (–1, 0) and (–4, 0) y–intercept: (0, 8) D. x–intercepts: (–1, 0) and (–4, 0) y–intercept: (0, 4)

OpenStudy (anonymous):

y intercept mean x =0 y(0) = 8 (0,8) x intercept : 0 = 2x^2 -10x + 8 2x^2 -8x-2x+8 2x(x-4)-2(x-4) (2x-2)(x-4) x = 1 , x = 4 (1,0) , (4,0) A

OpenStudy (anonymous):

You're a lifesaver <3 Thank you so much.